angle x and angle y are complementary angles. If angle x = 3x+7 and angle y =6x+20, find the value of x and the degree measure of each angle.
Answer by cleomenius(959)

3x + 7 + 6x + 20 = 90
9x = 63
x = 7
3(7) + 7 = 28 degrees.
6(7) + 20 = 62 degrees.
28 degrees + 62 degrees = 90 degrees.
This does check.
Cleomenius.